OpenStudy (anonymous):

How do I convert f(x) = 4 sin(2x - π) + 3 into a cosine function? Help please!

OpenStudy (anonymous):

In the first part I was supposed to identify similarities and differences. But based on those I'm still unsure of how to do that.

OpenStudy (freckles):

have you thought about co-function identity

OpenStudy (freckles):

\[\sin(u)=\cos(\frac{\pi}{2}-u)\]

OpenStudy (anonymous):

Oh! Well that helps a lot lol

OpenStudy (anonymous):

I think I can solve it from here but if I need anymore help I'll tag you :D

OpenStudy (freckles):

k